Physics 8lB Lab 8- Diffraction rev 4.0 Part 4: Diffraction Gratings Diffraction gratings are slides made up of a large number of very tiny slits. Although they bare the name "diffraction gratings", the underlying principle is again the superposition of the waves that pass through each slit. In this case the slits are usually taken to be small enough to approximate point sources, but we still have many slits to contend with, The result is that we end up with very bright and sharp (ie. narrow) intensity maxima at certain positions, and (nearly) complete cancellation at all the other positions Diff. Grating-Intensity Maxima: dsin@w ,-λ,0,A2A,3A".. (Slit spacing-d)
b) Find an expression relating ted and 0red (to the first maximum) to the values if you Does your expression used a green laser λgreen and θgreen (to the first maximum. depend on the slit spacing of the diffraction grating?
Answer #1

For the diffraction grating the Maxima are given at points,

d sin \theta = n \lambda

For the first maxima for red we get,

d sin (\theta_{red}) = \lambda_{red}

For the first maxima for green we get,

d sin (\theta_{green}) = \lambda_{green}

Dividing both the equations we get,

\Rightarrow \frac{sin(\theta_{red})}{sin (\theta_{green})} = \frac{\lambda_{red}}{\lambda_{green}}

Hence the expression relating the quantities does not depend on the slit spacing of diffraction grating.